OpenStudy (anonymous):

The difference of the areas of two squares is 315 square feet. Each side of the larger square is six times the length of a side of the smaller square. Find the length of a side of each square.

OpenStudy (anonymous):

let x be the larger square. The sides are 6y. ---- | | 6y ---- That is x=6y We know x^2 - y^2 = 315

OpenStudy (anonymous):

You have two equations and 2 unknowns. The answer is pretty. Let me know if you need help with the algebra!

OpenStudy (anonymous):

x^2 - 315 = (x/6)^2 and so on.

OpenStudy (anonymous):

Did you get it?